A Bibliography of Publications in ACM SIGPLAN Notices, 1960–1969

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 3.12

Title word cross-reference

+ [19].

1.5 [45].

68 [29, 21].

ALGOL [3, 8, 13, 15, 16, 19, 21, 29, 18]. Algol-based [18]. ALGOL-N [19]. algorithm [22]. Alternatives [38]. aspects [30]. authors [35].

base [30]. based [18]. Basel [30]. Bounce [23]. Bounce-and-skip [23]. Bulletin [41, 3, 8, 13, 15, 16, 10, 12, 2, 5, 9, 1, 4, 7, 14].

collateral [18]. Commenting [46].

Comments [25]. compile [36]. compiler [40]. concept [39]. considered [50]. context [22]. context-free [22]. control [23]. conversion [25].

DDI [24]. defining [18]. directing [23]. Document [47].

EPS [34]. evaluation [24]. execution [18]. extensibility [36, 39]. extensible [37, 30, 29, 38, 27, 40]. extension [33, 31].

facilities [31]. facility [33, 30]. feature [42]. features [32]. flow [23]. FORTRAN [12, 10]. free [22].

general [28, 18]. generation [21]. Go [50, 51]. GPL [28]. harmful [50]. hybrid [18].

IMP [31]. infinite [20, 21]. information [10, 12, 2]. interpreter [26]. Introduction [27]. involving [21].

JOVIAL [24, 2].

laboratory [44]. language [28, 30, 29, 18, 35, 32]. Languages [17, 37, 22, 11, 38, 27, 40]. LISP [41, 45, 42, 47, 48, 26, 49, 46]. Literate [52].

machine [22]. machine-oriented [22]. Macro [34]. Manual [43]. MLISP [43]. modes [20, 21]. Motivation [37].

N [19]. NELIAC [6]. news [6]. No [41, 6, 3, 8, 13, 15, 16, 10, 12, 2, 5, 9, 1, 4, 7, 14]. Number [49].

oriented [22].

Package [49]. Panel [35, 39]. particular [18]. PL [25, 36, 5, 9]. PL/I [25, 36, 5, 9]. polymorphic [32]. PPL [32]. precision [25]. procedures [18]. processing [34]. Program [47]. Programming [51, 52, 17, 11, 18, 32, 44]. Programs [46, 21, 23]. proposed [25]. Proteus [33]. purpose [28].

Random [49]. recognition [22]. Review [24, 17]. Revision [25]. rules [25].

seminar [17]. semiotic [11]. skip [23]. SNOBOL [4, 7, 14, 1]. Snobol4 [26]. Some [30, 40, 32]. Standard [48]. statement [50]. Statements [51]. Structure [47]. Structured [51]. study [24]. Systems [45].

technique [23]. techniques [40]. theory [11]. time [36]. Today [17]. Transformations [33].

UCC [17]. use [40]. User [43].

References

Waite:1967:SBNa

[1] William M. Waite. Snobol bulletin no. 1. ACM SIGPLAN Notices, 2(5):1–4, May 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

                                            org/doi/10.1145/1139522.1139528.

Oppenheim:1967:JIB

[2] David K. Oppenheim. Jovial information bulletin no. 1. ACM SIG-PLAN Notices, 2(5):1–5, May 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

                                            (electronic). URL https://dl.acm.org/
                                            doi/10.1145/1139522.1139524.

Duncan:1967:ABNa

[3] F. G. Duncan. ALGOL bulletin no. 25. ACM SIGPLAN Notices, 2(5):1–30, May 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

                                            org/doi/10.1145/1139522.1139526.

Waite:1967:SBNb

[4] William M. Waite. SNOBOL bulletin no. 2. ACM SIGPLAN Notices, 2(7):1–7, July 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

                                            org/doi/10.1145/1139514.1139516.

Southworth:1967:PBNa

[5] R. N. Southworth. PL/I bulletin no. 4. ACM SIGPLAN Notices, 2(8):1–40, August 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/doi/10.1145/1139518.1139520.

Anonymous:1967:NNN

[6] Anonymous. NELIAC news no. 1. ACM SIGPLAN Notices, 2(9):1–6, September 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/doi/10.1145/1139530.1139532.

Waite:1967:SBNc

[7] William M. Waite. SNOBOL bulletin no. 3. ACM SIGPLAN Notices, 2(10):1–7, October 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/1139494.
1139496.

Duncan:1967:ABNb

[8] F. G. Duncan. ALGOL bulletin no. 26. ACM SIGPLAN Notices, 2(11):1–49, November 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/1139498.
1139500.

Southworth:1967:PBNb

[9] R. N. Southworth. PL/I bulletin no. 5. ACM SIGPLAN Notices, 2(12):1–71, December 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/1139502.
1139504.

Golde:1968:FIBa

[10] Hellmut Golde. Fortran information bulletin no. 1. ACM SIGPLAN Notices, 3(2):1–5, February 1968. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL https://dl.acm.org/doi/
                                        10.1145/1139544.1139546.

Laszlo:1968:TST

[11] Steven I. Laszlo. Toward a semiotic theory of programming languages. ACM SIGPLAN Notices, 3(5):24–27, May 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/954348.954349.

Golde:1968:FIBb

[12] Hellmut Golde. FORTRAN information bulletin no. 3. ACM SIGPLAN Notices, 3(7–8):1–19, July–August 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duncan:1968:ABNa

[13] F. G. Duncan. ALGOL bulletin no. 27. ACM SIGPLAN Notices, 3 (9):1–65, September 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL https://dl.acm.org/doi/
                                        10.1145/1139560.1139562.

Waite:1968:SBN

[14] W. M. Waite. SNOBOL bulletin no. 5. ACM SIGPLAN Notices, 3(10):1–6, October 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/1139552.
                                        1139554.

Duncan:1968:ABNb

[15] F. G. Duncan. ALGOL bulletin no. 28. ACM SIGPLAN Notices, 3(11):1–83, November 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/1139536.
1139538.

Duncan:1968:ABNc

[16] F. G. Duncan. ALGOL bulletin no. 29.

ACM SIGPLAN Notices, 3(12):1–69, December 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/1139540.
1139542.

Shaw:1969:RUS

[17] C. J. Shaw. Review of UCC seminar on “Programming Languages Today”. ACM SIGPLAN Notices, 4(2):24–25, February 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/doi/10.1145/953392.953393.

Roberts:1969:ABL

[18] John B. Roberts. An Algol-based language for defining collateral execution of procedures in general and hybrid programming in particular. ACM SIGPLAN Notices, 4(3):17–37, March 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/doi/10.1145/1316865.1316866.

Igarashi:1969:A

[19] S. Igarashi, T. Iwamura, K. Sakuma, T. Simauti, T. Simuzu, S. Takasu, E. Wada, and N. Yoneda. ALGOL-N^{+}. ACM SIGPLAN Notices, 4(3):38–85, March 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/1316865.
1316867.

Koster:1969:IM

[20] C. H. A. Koster. On infinite modes. ACM SIGPLAN Notices, 4(3):86–89, March 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/1316865.1316868.

Meertens:1969:GAP

[21] L. Meertens. On the generation of ALGOL 68 programs involving infinite modes. ACM SIGPLAN Notices, 4(3):90– 92, March 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/1316865.
                                        1316869.

Floyd:1969:MOR

[22] Robert W. Floyd. A machine-oriented recognition algorithm for context-free languages. ACM SIGPLAN Notices, 4(5):28–29, May 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL https://dl.acm.org/doi/
                                        10.1145/987470.987472.

Nievergelt:1969:BST

[23] J. Nievergelt and M. I. Irland. Bounceand-skip: a technique for directing the flow of control in programs. ACM SIG-PLAN Notices, 4(5):29–31, May 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

                                        (electronic). URL https://dl.acm.org/
                                        doi/10.1145/987470.987473.

Shaw:1969:RDJ

[24] Christopher J. Shaw. Review of the DDI JOVIAL evaluation study. ACM SIG-PLAN Notices, 4(6):13–17, June 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/
doi/10.1145/1317443.1317444.

Weinberg:1969:CPR

[25] G. M. Weinberg. Comments on the proposed “Revision of PL/I conversion and precision rules”. ACM SIG-PLAN Notices, 4(6):17–20, June 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/
doi/10.1145/1317443.1317445.

McGowan:1969:LIS

[26] Clement McGowan and Peter Wegner. A LISP interpreter in Snobol4. ACM SIG-PLAN Notices, 4(6):20–33, June 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/
doi/10.1145/1317443.1317446.

Perlis:1969:IEL

[27] Alan J. Perlis. Introduction to extensible languages. ACM SIGPLAN Notices, 4(8):3–5, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115859.

Garwick:1969:GGP

[28] Jan V. Garwick. GPL, a general purpose language. ACM SIGPLAN Notices, 4(8):6–8, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115860.

Mailloux:1969:AEL

[29] Barry J. Mailloux and John E. L. Peck.

Algol 68 as an extensible language. ACM SIGPLAN Notices, 4(8):9–13, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/doi/10.1145/1115858.1115861.

Jorrand:1969:SAB

[30] Philippe Jorrand. Some aspects of Basel, the base language for an extensible language facility. ACM SIG-PLAN Notices, 4(8):14–17, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/
doi/10.1145/1115858.1115862.

Irons:1969:EFI

[31] Edgar T. Irons. The extension facilities of IMP. ACM SIGPLAN Notices, 4(8):18–19, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115863.

Standish:1969:SFP

[32] Thomas A. Standish. Some features of PPL, a polymorphic programming language. ACM SIGPLAN Notices, 4(8):20–26, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115864.

Bell:1969:TEF

[33] James R. Bell. Transformations: the extension facility of Proteus. ACM SIGPLAN Notices, 4(8):27–31, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/doi/10.1145/1115858.1115865.

MacLaren:1969:MPE

[34] M. Donald MacLaren. Macro processing in EPS. ACM SIGPLAN Notices, 4(8):32–36, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115866.

Shaw:1969:PLA

[35] Christopher J. Shaw. Panel of language authors. ACM SIGPLAN Notices, 4(8):37–39, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115867.

Nicholls:1969:PCT

[36] John Nicholls. PL/I compile time extensibility. ACM SIGPLAN Notices, 4(8):40–44, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115868.

Cheatham:1969:MEL

[37] Thomas E. Cheatham. Motivation for extensible languages. ACM SIG-PLAN Notices, 4(8):45–49, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/
doi/10.1145/1115858.1115869.

McIlroy:1969:AEL

[38] M. D. McIlroy. Alternatives to extensible languages. ACM SIGPLAN Notices, 4(8):50–52, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115870.

Wegner:1969:PCE

[39] Peter Wegner. Panel on the concept of extensibility. ACM SIG-PLAN Notices, 4(8):53–54, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160

(electronic). URL https://dl.acm.org/
doi/10.1145/1115858.1115871.

Standish:1969:SCC

[40] Thomas A. Standish. Some compiler– compiler techniques for use in extensible languages. ACM SIGPLAN Notices, 4(8):55–62, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1115858.1115872.

Bobrow:1969:LB

[41] Daniel G. Bobrow. LISP bulletin no. 1. ACM SIGPLAN Notices, 4 (9):17–57, September 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/doi/
10.1145/1132291.1132032.

Bobrow:1969:NLF

[42] Daniel G. Bobrow. A new LISP feature. ACM SIGPLAN Notices, 4(9): 17, September 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Smith:1969:MUM

[43] David Canfield Smith. MLISP user’s manual. ACM SIGPLAN Notices, 4 (9):18, September 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Teitelman:1969:TPL

[44] Warren Teitelman. Toward a programming laboratory. ACM SIGPLAN Notices, 4(9):18, September 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Bobrow:1969:LS

[45] Daniel G. Bobrow. LISP 1.5 systems. ACM SIGPLAN Notices, 4(9): 19, September 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Shaw:1969:CLP

[46] Christopher J. Shaw. Commenting LISP programs. ACM SIGPLAN Notices, 4 (9):20, September 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Bobrow:1969:PDL

[47] Daniel G. Bobrow. A program to document LISP program structure. ACM SIGPLAN Notices, 4(9):21–24, September 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Hearn:1969:SL

[48] Anthony C. Hearn. Standard LISP. ACM SIGPLAN Notices, 4(9):25–52, September 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Palme:1969:LRN

[49] Jacob Palme. A LISP random number package. ACM SIGPLAN Notices, 4(9): 53–57, September 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Embedded in [41].

Dijkstra:1968:GSC

[50] Edsger Wybe Dijkstra. Go to statement considered harmful. Communications of the ACM, 11(3):147–148, March 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This paper inspired scores of others, published mainly in SIGPLAN Notices up to the mid-1980s. The best-known is [51].

Knuth:1974:SPG

[51] Donald E. Knuth. Structured programming with go to statements. ACM Computing Surveys, 6(4):261–301, December 1974. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). Reprinted with revisions in Current Trends in Programming Methodology, Raymond T. Yeh, ed., 1 (Englewood Cliffs, NJ: Prentice-Hall, 1977), 140–194; Classics in Software Engineering, Edward Nash Yourdon, ed. (New York: Yourdon Press, 1979), 259–321. Reprinted with “final” revisions in [52, pp. 17–89]. This paper is a response to [50].

Knuth:1992:LP

[52] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xvi + 368 pp. LCCN QA76.6 .K644 1992.

US$24.95 (paperback), US$59.95 (hardcover). Distributed by the University of Chicago Press.